OpenStudy (anonymous):

Terri’s Ice Cream Parlor makes groumet ice cream cones. The volume of each cone is 8(pi) cubic inches. The height is 4 more inches than the radius of the cone’s opening. Write a polynomial equation that represents the volume of an ice cream cone. Use the formula for the volume of a cone, v=1/3(pi)r^2h

OpenStudy (anonymous):

Okay thank you but now it is asking for the possible values of r

OpenStudy (anonymous):

V = (π/3) r² h = 8 π r² h = 24 π r² (r + 4) = 24 π( r³ + 4r² ) - 24 = 0

OpenStudy (anonymous):

That's why you need to solve the equation :)

OpenStudy (anonymous):

How do you do that? I'm sorry I really don't understand polynomials

OpenStudy (anonymous):

Nevermind I got it! Thankyou